OpenStudy (anonymous):

Paper plates come in packages of 15 or 20. George bought 10 packages and had a total of 170 plates. How many of each package did he buy? 3 packages of 15, 7 packages of 20 7 packages of 15, 3 packages of 20 4 packages of 15, 6 packages of 20 6 packages of 15, 4 packages of 20

OpenStudy (pfenn1):

Let x = number of 15-plate packages bought, then the number of 20-plate packages = 10 - x. The equation to find x would be\[15x+20(10-x)=170\] Solve for x.

OpenStudy (anonymous):

x = 1?

OpenStudy (pfenn1):

No. \[15x+200-20x=170\]What is x?

OpenStudy (anonymous):

x = 6

OpenStudy (anonymous):

is it D?

OpenStudy (pfenn1):

Yes, x=6. Sorry. I was helping someone else.

OpenStudy (anonymous):

so it was D right?

OpenStudy (anonymous):

or C?

OpenStudy (anonymous):

you can model this situation by solving the system: let x = # 15plate pkg let y = # 20 plate pkg x+y=10 15x + 20y = 170 best way i think is to just check against your choices..

OpenStudy (pfenn1):

I defined x above. You got x=6 for an answer. So which is correct C or D?